# Generalized Simple Graphical Rules for Assessing Selection Bias

### v1

@@ 6 Discussion
Our work assesses and extends the utility of existing simple graphical rules for addressing selection bias. Derived based on the fundamental principle of reading conditional independence based on d-separation, existing simple graphical rules are inherently coupled with certain identification formulas and applicable to resolving selection bias for a specific subset of selection mechanisms. We show that, some important existing rules [6, 7, 8] can have stringent requirements on covariates adjustment to avoid confounding bias while correcting selection bias. For example, the adjustment for post-treatment variables is generally discouraged by these rules even if it helps separate the selection from the outcome. It is worth noting that while post-treatment variables should not be adjusted for in addressing confounding bias, these variables can indeed be accounted for when dealing with selection bias. We use two concrete examples to demonstrate the limitations of existing rules and their identification strategies in addressing selection bias.
@@
Specific to the two examples, we theoretically prove the identifiability of the ATE by the g-formula and the IPW method when relevant external information in the full population is available. The g-formula relies on adjusting for the post-treatment upstream cause of selection using a conditional distribution, while the IPW method seeks to balance the distributions of variables by modeling the probability of selection in the full population. The characteristics of these approaches are related to different sufficient conditions that we propose for identifying the ATE. They can be applied to complex observational settings with observed confounders for the relationships between the treatment and post-treatment variables, and unobserved confounders for the relationships between the upstream cause of selection and the outcome. For the IPW method, we also propose two estimators and their variance estimators for the statistical inference of the ATE. We outline the procedures and conduct simulation analysis to verify the performance of our estimators, which correct the serious bias by a complete-case analysis using the selected sample.
@@
Beyond the two examples, our approach to correcting the selection bias validates and suggests possible directions for identifying the ATE when conventional approaches fail. First, researchers have shown that the g-formula and the IPW method are reliable alternatives to the oftentimes biased selected-sample analysis, for addressing selection bias and beyond [25, 14, 17, 18, 19, 11, 20, 21, 22]. Second, it is possible to adjust for post-treatment variables to recover the intended treatment effect without incurring overadjustment bias [26, 27, 28]. The adjustment is in a broader sense that its conditional distribution might be marginalized, or it might be included in a model for the selection, compared to a simple inclusion of it in the outcome regression. For certain data dependence structures and research questions, principal stratification based on post-treatment variables is a powerful framework for identifying numerous treatment effects of interest [29]. Finally, the availability of external data beyond the selected sample in the entire target population could be extremely useful in inviting novel identification strategies. The identification strategies are closely related to and dependent on the data collection process and the associated assumptions. Ad-hoc analyses may be necessary to address various data challenges based on fundamental graphical rules [6].
@@ Acknowledgements
We are grateful to Forrest Crawford, Laura Forastiere, Fan Li, and Thomas Thornhill for helpful comments.

### v3

@@ 6 Discussion
Our work assesses and extends the utility of existing graphical rules for addressing selection bias. Derived based on the fundamental principle of reading conditional independence using d-separation, existing simple graphical rules are inherently coupled with certain identification formulas and applicable to resolving selection bias for a specific subset of selection mechanisms. We show that, some important existing rules [6, 7, 8] can have stringent requirements on covariate adjustment to avoid confounding bias while correcting selection bias. For example, the adjustment for post-treatment variables is generally discouraged by these rules even if it helps separate the selection from the outcome. It is worth noting that while post-treatment variables should not be adjusted for in addressing confounding bias, these variables can indeed be accounted for when dealing with selection bias. We use two concrete examples using causal diagrams to demonstrate the limitations of existing rules and their identification strategies in addressing selection bias.
@@
Specific to the two examples, we theoretically prove the identifiability of the ATE by the g-formula and the IPW method when relevant external information in the full population is available. The g-formula relies on adjusting for the post-treatment upstream cause of selection using a conditional distribution, while the IPW method seeks to balance the distributions of variables by modeling the probability of selection in the full population. The characteristics of these approaches are related to different sufficient conditions that we propose for recovering the ATE. They can be applied to complex observational settings with observed confounders for the relationships between the treatment and post-treatment variables, and unobserved confounders for the relationships between the upstream cause of selection and the outcome. For the IPW method, we also propose two estimators and their variance estimators for the statistical inference of the ATE. We outline the procedures and conducte simulation analysis to verify the performance of our estimators, which correct the serious bias by a complete-case analysis using the selected sample.
@@
Beyond the two examples, our approach to correcting the selection bias validates and suggests possible directions for recovering the ATE when conventional approaches fail. First, researchers have shown that the g-formula and the IPW method are reliable alternatives to the oftentimes biased selected-sample analysis, for addressing selection bias and beyond [25, 14, 17, 18, 19, 11, 20, 21, 22]. Second, it is possible to adjust for post-treatment variables to recover the intended treatment effect without incurring overadjustment bias [26, 27, 28]. The adjustment is in a broader sense that its conditional distribution might be marginalized, or it might be included in a model for the selection, rather than a simple inclusion of it in the outcome regression.
@@ Acknowledgements
This section has been temporarily removed from the manuscript for peer review.